OpenStudy (anonymous):

f(X) =2 (√5-x) g(x)= (√x^2-9) find f+g how do i combine these

OpenStudy (anonymous):

put a plus sign between them. really that is it

OpenStudy (anonymous):

oh wow. so u cant add the square roots together right? thank you

OpenStudy (anonymous):

you cannot combine then in to one radical though!

OpenStudy (anonymous):

i am assuming it is \[f(x)=2\sqrt{5-x}\] \[g(x)=\sqrt{x^2-9}\] so \[(f+g)(x)=2\sqrt{5-x}+\sqrt{x^2-9}\]

OpenStudy (anonymous):

thanks